OpenStudy (anonymous):

Consider the angle that measures 5pi/4 a) determine TWO coterminal angles (one positive and one negative) b)Convert the angle to degree measure.

OpenStudy (anonymous):

\[\huge \frac{5 \pi}{4} =\frac{5 \pi}{4} \times \frac{180^0}{\pi} = 225^0\] PositiveCoterminal angles= \[\huge 360^0 +225^0 =585^0\] Negative coterminal angle= \[\huge 225^0 -360^0 =-135^0\] @Klong0091

OpenStudy (anonymous):

Thank you so much !!!!!!!!!

OpenStudy (anonymous):

@Klong0091 Do I deserve a Medal for this ?

OpenStudy (anonymous):

yeah thanks